OpenStudy (anonymous):

Estimate by first rounding each number to the nearest 1/2 . -3 8/9 x 1 1/2 A. –6 B. –3 C. 6 D. 8 Is 8 correct?

OpenStudy (dangerousjesse):

How did you get 8? :)

OpenStudy (anonymous):

-3 8/9 = -4 1 1/2= 1 @DangerousJesse tbh it was a guess, what I really got was -4, although it isn't one of the answer choices :c

OpenStudy (anonymous):

@DangerousJesse Are you going to help?

OpenStudy (anonymous):

oh ok thanks

pooja195 (pooja195):

It's right

OpenStudy (dangerousjesse):

Oh, I'm so sorry, I got distracted.

OpenStudy (anonymous):

@DangerousJesse It's ok :3

OpenStudy (anonymous):

The prompt asked you to first round to the nearest *half*. \[-3 \frac{ 8 }{ 9 } ~~\rightarrow~ -4\]\[ ~~~1 \frac{ 1 }{ 2 } ~~\rightarrow~~~~ 1 \frac{ 1 }{ 2 }\] And if we multiply these values, we would get this product: \[-4 \times 1.5 = -6\]

OpenStudy (anonymous):

For anyone looking for the right answer, it's -6 like Cardinal said (I took the assesment for k12 and got 80%) I looked at the wrong answers which was this one :)